Find (z)*2020/3 if z=cos11π/6 +isin7π/6.


1
Expert's answer
2021-08-02T16:08:12-0400

We know that


"\cos \dfrac{11\pi}{6}=\cos \dfrac{\pi}{6}=\dfrac{\sqrt{3}}{2}"

"\sin \dfrac{7\pi}{6}=-\sin \dfrac{\pi}{6}=-\dfrac12"

So, then


"z=\dfrac{\sqrt{3}}{2}-\dfrac12i"

Finally,


"z\cdot \dfrac{2020}{3}=\dfrac{1010\sqrt{3}}{3}-\dfrac{1010}{3}i."